Troubleshooting
Error number/message |
See also
Choose a keyword to go to troubleshooting or other technical information associated with it.
- 10061 or 10062: Lid leaks on Morgan cabinets
- 10063: External
debris blower system blocked (high vacuum) (VLF devices only)
- 10064: External debris blower not connected or filters full
- 19002: Debris removal vacuum level low (ThermoFlex devices only)
- 27002/43002: Particulate only collection (POC) debris removal is not enabled on this device
- 27003/43003: Particulate and chemical collection (PACC) debris removal is not enabled on this device
- 27004/43004: The air jet is enabled, but not turned on
- 27005/43005: The blower is not turned on
- 27006/43006: The flow rate is below the minimum required
- 27007/43007: Cabinet has lost power
- 27008/43008: The particulate filter is not present or upside down, the lid is not closed, or data cable is disconnected
- 27009/43009: The particulate filter is almost full. Please ensure a replacement filter is on hand
- 27010/43010: The particulate filter is full. You must replace the particulate filter to continue imaging
- 27011/43011: Debris removal collection nozzle is blocked
- 27012/43012: The debris hose is disconnected
- 27013/43013: There is a problem with one or more of the pressure sensors in the debris removal cabinet
- 27014/43014: The debris hose is disconnected, the cabinet has lost power, or the cabinet lid is not latched properly
- Blower motor failures (error #27014/43014)
- 27015/43015: There may be a partial blockage of the collection nozzle, or leak, but flow is still above minimum
- 27016/43016: There is a partial blockage of the collection nozzle, or a leak and the flow rate is below minimum
- 27017/43017: The media being imaged releases potentially harmful emissions. See the Operator Manual for more information
- 27019/43019: The chemical filter has almost expired. Please ensure a replacement filter is on hand
- 27020/43020: The chemical filter has expired. You must replace the chemical filter and reset the output device to continue imaging
- 27021/43021: The chemical filter usage sensor is corrupt or the chemical filter was changed and the output device was not reset
- 27022/43022: Unable to communicate with the chemical filter
- 27023/43023: The chemical filter usage calculated is negative
- 27024/43024: Unable to calibrate pressure sensor. Where = nozzle, bellmouth, or differential as appropriate
- 27025/43025: Unable to communicate with the chemical filter during initialization
- 27026/43026: Timeout: the pressure sensor output failed to stabilize
- 27027/43027: No debris removal options are enabled on this device
- 27028/43028: There is a problem with the nozzle pressure sensor in the output device
- 27029/43029: Command failed: cannot run more than one PACC option at a time
- 27030/43030: Cannot start new debris option while another is in an error condition
- 27031/43031: The chemical filter installed is not suitable for the selected media
- 27032/43032: The particulate filter is almost full AND the chemical filter has almost expired
- 34039: Restricted parameter, insufficient access level